OpenStudy (anonymous):

How can I solve: sin(ln(x+1000)-sin(ln(x)) when x approaches infinity

OpenStudy (anonymous):

Use the Mean Value Theorem for \[ g(x)=\sin (\ln x) \text { in the interval } [x, 1000+x] \]

OpenStudy (anonymous):

There is a \(c_x\ \in [x,1000+x]\) such that \[ g(x+1000)-g(x) = 1000 g'(c_x) \\ |\sin (\ln( x + 1000) -\sin(\ln x)| =| 1000 \ \frac{\cos(\ln( c_x)} {c_x}|\le \\ 1000 \frac 1 {x_c}\le 1000 \frac 1 {x} \] Can you finish it now?

OpenStudy (anonymous):

Sure!! Your response is awesome. Now I know how to deal with the problem.

OpenStudy (anonymous):

YW